OpenStudy (anonymous):

Help please? A certain business has implemented a new incentive plan whereby an employee earns tokens for every sale they make. For 1 sale, an employee earns 1 token. For 2 sales, an employee earns 8 tokens. For 3 sales, an employee earns 17 tokens. For 4 sales, an employee earns 28 tokens. How many tokens will be earned for eight sales? 92 97 104 113

OpenStudy (anonymous):

@welkerlover1283

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Just a sec, almost have it

OpenStudy (anonymous):

Well I assume this a pattern problem so just find difference between each sale increments 4 sales: 28 tokens -3 sales: 17 tokens _____________________________ X 3 sales: 17 tokens -2 sales: 8 tokens _____________________________ X 2 sales: 8 tokens -1 sales: 1 tokens _____________________________ X Then you find a pattern

OpenStudy (anonymous):

How do I find a pattern?

OpenStudy (anonymous):

Minus the tokens from 4sales and 3sales then 3sales and 2sales then 2sales and 1sale.

OpenStudy (ranga):

1, 8, 17, 28, ...... ? You will notice 7 is added to the first to get the second 9 is added to the second to get the third 11 is added to the third to get the fourth What will be added to the fourth to get the fifth?

OpenStudy (anonymous):

41 and 56?

OpenStudy (anonymous):

Oh! Is the answer 113?

OpenStudy (ranga):

Yes. 41 and 56 are your fifth and sixth. What is the seventh and eighth term?

OpenStudy (anonymous):

113 :)

OpenStudy (anonymous):

is the 8th

OpenStudy (anonymous):

7th term is 92 but that doesn't matter haha the answer is 113 right?

OpenStudy (ranga):

No. 7th term is 56 + 17 = 73. 8th term is 73 + 19 = 92

OpenStudy (anonymous):

ohhh lol

OpenStudy (ranga):

Add: 7 9 11 13 15 17 19 1 8 17 28 41 56 73 92

OpenStudy (ranga):

1 + 7 = 8 8 + 9 = 17 17 + 11 = 28 and so on and so forth.

OpenStudy (anonymous):

thank you :)

OpenStudy (ranga):

you are welcome.